OpenStudy (izzy18970):

Solve this proportion. 1/10+1/7=1/x

OpenStudy (izzy18970):

I need help with the steps

OpenStudy (izzy18970):

no

OpenStudy (anonymous):

add on the left, then flip it

OpenStudy (anonymous):

what is \[\frac{1}{7}+\frac{1}{10}\]?

OpenStudy (izzy18970):

you have to find the least common denominator right?

OpenStudy (izzy18970):

70

OpenStudy (anonymous):

i guess but here it is just 70 for the lcd

alones (alones):

17/70=1/x \ 1: Cross-multiply. 17/70=1/x 17*x=(1)*(70) 17x=70 2: Divide both sides by 17. 17x/17=70/17 x=70/17 @satellite73 did i solve this right?? ;3

OpenStudy (anonymous):

\[\frac{1}{a}+\frac{1}{b}=\frac{a+b}{a\times b}\]

OpenStudy (anonymous):

@AloneS yes you did , except for the words "cross multiply" which do not apply

alones (alones):

Oh my bad xP Thank you for the correction<3

OpenStudy (izzy18970):

so the answer is 4.1 ?

OpenStudy (izzy18970):

4.12

OpenStudy (anonymous):

the answer is not a decimal, it is fraction

OpenStudy (izzy18970):

oh ok thank you

OpenStudy (anonymous):

\[\frac{1}{7}+\frac{1}{10}=\frac{10+7}{10\times 7}=\frac{17}{70}\] if \[\frac{17}{70}=\frac{1}{x}\] then \[\frac{70}{17}=x\]